OpenStudy (anonymous):

3x+2y=11 and x+y=4 substitution method ?

OpenStudy (aripotta):

what do you get when you solve for x in the second equation?

OpenStudy (anonymous):

idk that all the problem gave me

OpenStudy (aripotta):

yea, you have to solve for a variable first before you can substitute anything. so solve for x in terms of y in the second equation

OpenStudy (aripotta):

x+y=4 just rearrange the equation so that it's just x on the left side

OpenStudy (anonymous):

yea i get what your saying but my issue is getting the hang of how to set these equations up on my own

OpenStudy (aripotta):

do you have more questions like this?

OpenStudy (anonymous):

yea i had an equation similar to this one but i got that one answered it wass 5x+3y=7 and y= x-3

OpenStudy (aripotta):

i can do this one for ye; show you all the steps and stuff. you can ask me questions if you have any. then you'll try that next one, mmk?

OpenStudy (anonymous):

What you can do is use one of the equations to express x in terms of y, and then substitute that into the other equation.

OpenStudy (aripotta):